# MK-SGN: A Spiking Graph Convolutional Network with Multimodal Fusion and Knowledge Distillation for Skeleton-based Action Recognition

### v4

@@ 4.2.2 The Effect of Different Components
Compared to Base-SGN*, MK-SGN delivers substantial accuracy gains of 14.3% and 11.3% on XSub and XView, respectively. Additionally, when compared to the classic ANN-based model ST-GCN, MK-SGN achieves a remarkable 94.78% reduction in energy consumption while maintaining comparable accuracy. These results highlight the efficiency and effectiveness of MK-SGN, showcasing its suitability for energy-constrained applications without compromising performance.
4.3 Comparison with State-of-the-Art Methods
In this section, we conduct a comprehensive comparison between our proposed MK-SGN and SOTA methods for skeleton-based action recognition. The comparison is divided into two parts:(1). model architecture comparison and (2).metric comparison.
In the model architecture comparison, we focus on the design and layer configurations of MK-SGN in contrast to GCN and SNN models, using the NTU-RGB+D 60 dataset as a case study. This includes both publicly available results and those we reproduced under consistent experimental settings. The architecture analysis establishes the foundation for understanding the computational and operational differences, which are crucial for fair metric comparisons.
In the metric comparison, we evaluate MK-SGN against SOTA models across several dimensions, including accuracy, number of operations (OPs), and theoretical energy consumption. These metrics provide a quantitative perspective on MK-SGN’s ability to balance high performance with energy efficiency, particularly in energy-constrained environments.
Table 8: Model’s architecture for NTU-RGB+D 60
@@ 4.3 Comparison with State-of-the-Art Methods
final 256/class number GAP-FC 256/class number 256/class number GAP-SN-FC 256/class number GAP-FC
@@ 4.3.1 Model Architecture Comparison
Table 8 summarizes the architectural differences between our proposed MK-SGN and various SOTA models, including GCN, and Spikformer, as applied to the NTU-RGB+D 60 dataset. This comparison highlights how each model is structured to handle skeleton-based action recognition tasks, with adaptations made to ensure fair performance evaluations.
The GCN model serves as a classical ANN-based approach, employing GC layers combined with TC layers for spatio-temporal feature extraction. Its structure consists of 10 layers of GC-TC operations, with progressively increasing feature dimensions and downsampling via strides, culminating in a GAP-FC layer for classification. This architecture represents a strong baseline for ANN-based recognition.
The Base-SGN* model acts as the baseline SNN structure, leveraging SGC and STC layers to process spike-form data. It follows a similar configuration to GCN but is adapted for spiking computation. However, it lacks the advanced features introduced in MK-SGN.
The Spikformer model, a spiking-based Transformer, incorporates SSA layers and MLP components for feature extraction. For this study, we modified Spikformer to align with MK-SGN in terms of layer count and hidden dimensions, ensuring a fair comparison of metrics such as energy efficiency and accuracy. This adjustment enables us to evaluate the performance of Spikformer in a setting directly comparable to MK-SGN. Similarly, other network variations like Spiking-driven-former, Spiking-driven-former v2 and Spiking Wavelet Transformer used in the study are adapted to maintain consistency across evaluations.
@@ 4.3.2 Metric Comparison
Table 9: Comparative results on NTU RGB+D 60 and 120. We evaluate our model in terms of classification accuracy (%). X-Sub and X-View represent cross-subject and cross-view splits.
@@
Param OPs Power Csub Cview Csub CSet
(M) (G) (mJ) (60) (60) (120) (120)
@@
MK-SGN (Ours) SNN SGN 4 2.17 0.68 0.614 78.5 85.6 67.8 69.5
@@
We compare our proposed MK-SGN with the current SOTA ANNs and SNNs in skeleton-based action recognition, focusing on two critical performance metrics: OPs and theoretical energy consumption. The results are summarized in Table 9 and Table 10.
To ensure fairness, the reported results for existing multi-modal GCN methods utilize all four modalities: joint, bone, joint motion, and bone motion. For these methods, the softmax scores of multiple streams are fused by summing the scores from each stream. The OPs and theoretical energy consumption for these multi-modal methods are calculated by multiplying the values for a single stream by the total number of modalities.
@@
Additionally, we reproduce the results of SOTA SNN models using only the joint modality to provide a direct and fair comparison with MK-SGN. These reproduced SNN models are carefully adapted to align with MK-SGN in terms of the number of layers and hidden dimensions. This ensures consistency in the experimental settings and allows for a robust evaluation of their efficiency and effectiveness.
Figure 5: Theoretical Energy Consumption Comparison
(a) Comparsion on NTU-RGB+D 60 Csub
(b) Comparsion on NTU-RGB+D 60 Cview
(c) Comparsion on NTU-RGB+D 120 Csub
(d) Comparsion on NTU-RGB+D 120 Cset
Figure 6: Comparsion of Accuracy with SOTA SNN
The comparison highlights that MK-SGN strikes a remarkable balance between accuracy and energy efficiency. Compared to existing methods, MK-SGN achieves significant reductions in theoretical energy consumption, emphasizing its suitability for energy-constrained scenarios, such as edge computing and real-time applications. By leveraging its spiking-driven architecture, MK-SGN demonstrates the ability to deliver competitive or superior performance while maintaining exceptional energy efficiency. Detailed results and analysis are provided in the following subsections.
Results on NTU-RGB+D 60 and 120
As shown in Table 9, MK-SGN demonstrates superior performance on both NTU-RGB+D 60 and NTU-RGB+D 120 datasets, showcasing its remarkable energy efficiency and competitive accuracy.
@@
On the NTU-RGB+D 60 and NTU-RGB+D 120 datasets, MK-SGN achieves a significant reduction in theoretical energy consumption, requiring only 0.614 mJ per sample. This represents a 96.16% decrease compared to ST-GCN (16.01 mJ), a 99.63% decrease compared to 2S-AGCN (164.68 mJ), and a 99.80% decrease compared to MS-G3D (311.09 mJ). Even when compared to lightweight models like Shift-GCN (46 mJ) and CTR-GCN (36.25 mJ), MK-SGN reduces energy consumption by 98.67% and 98.31%, respectively, highlighting its remarkable efficiency advantage across all ANN-based GCN architectures, as showed in Figure 5.
Beyond energy savings, MK-SGN surpasses SOTA SNN-based models in both accuracy and efficiency, as highlighted in Table 9 and Figure 6. For NTU-RGB+D 60, MK-SGN achieves classification accuracies of 78.5% (Csub) and 85.6% (Cview), outperforming competitive SNN models such as Spikformer (73.9% Csub, 80.1% Cview) and Spike-driven Transformer V2 (77.4% Csub, 83.6% Cview). Importantly, MK-SGN achieves this with significantly reduced energy consumption, requiring only 0.614 mJ per sample compared to Spikformer’s 2.17 mJ and Spike-driven Transformer V2’s 2.91 mJ. MK-SGN also demonstrates an efficient parameter design with 2.17M parameters, far less than Spike-driven Transformer V2’s 11.47M, highlighting its lightweight and efficient architecture.
On the NTU-RGB+D 120 dataset, MK-SGN continues to outperform SNN models, achieving accuracies of 67.8% (Csub) and 69.5% (CSet). In comparison, Spikformer achieves 61.7% (Csub) and 63.7% (CSet), while Spike-driven Transformer V2 achieves 64.3% (Csub) and 65.9% (CSet). Despite these improvements in accuracy, MK-SGN remains highly energy-efficient, consuming only 0.614 mJ, compared to Spiking Wavelet Transformer’s 2.01 mJ and Spike-driven Transformer V2’s 2.91 mJ. This demonstrates MK-SGN’s superior performance and efficiency across datasets.
These results validate MK-SGN as a robust and energy-efficient solution for skeleton-based action recognition, excelling in scenarios requiring both high accuracy and low computational cost. By leveraging its spiking-driven architecture, MK-SGN sets a new benchmark for balancing performance and efficiency in SNN.
Results on NW-UCLA
@@
On the NW-UCLA dataset, as summarized in Table 10, MK-SGN achieves a classification accuracy of 92.3%, surpassing SOTA SNN models. It outperforms models like Spikformer (85.4%) and Spike-driven Transformer V2 (89.4%) while maintaining significantly lower energy consumption, consuming only 0.207 mJ per sample compared to 0.914 mJ for Spike-driven Transformer V2.
This highlights MK-SGN’s ability to effectively balance accuracy and energy efficiency without compromising its computational performance. Its lightweight structure and optimized spiking modules allow for efficient recognition, particularly in scenarios where minimizing resource consumption is critical. These results further emphasize MK-SGN’s versatility and practicality across diverse application settings.
@@ 5 Conclusion
In this work, we propose the Spiking Graph Convolutional Network (MK-SGN), a novel architecture that combines the strengths of SNNs and graph convolutional networks GCNs for skeleton-based action recognition. This framework is, to the best of our knowledge, the first to effectively integrate spiking neural computation with graph-based modeling, offering a new perspective on energy-efficient and high-performance recognition methods.
The MK-SGN addresses the critical challenge of high energy consumption in traditional GCN-based methods while maintaining competitive performance. By leveraging multimodal fusion and knowledge distillation, MK-SGN not only optimizes the use of spiking neural computation but also enhances the ability to extract and integrate meaningful features across modalities. These innovations enable MK-SGN to significantly reduce energy consumption on three benchmark skeleton action recognition datasets, achieving over 98% reduction compared to the best ANN-based methods.
Moreover, MK-SGN consistently surpasses SOTA SNN models in recognition accuracy, achieving an average improvement of 5%. This highlights its ability to balance energy efficiency and computational effectiveness, making it particularly well-suited for real-time and resource-constrained applications. These results establish MK-SGN as a robust and practical solution for skeleton-based action recognition, paving the way for future exploration of spiking-based multimodal frameworks in other domains.

### v5

@@ 4.2.2 The Effect of Different Components
Compared to Base-SGN*, MK-SGN delivers substantial accuracy gains of 14.3% and 11.3% on XSub and XView, respectively. Additionally, when compared to the classic ANN-based model ST-GCN, MK-SGN achieves a remarkable 96.28% reduction in energy consumption while maintaining comparable accuracy. These results highlight the efficiency and effectiveness of MK-SGN, showcasing its suitability for energy-constrained applications without compromising performance.
Sensitivity Analysis Experiments
After evaluating the effect of different components in MK-SGN, we conduct sensitivity analysis on key model hyperparameters to assess their impact on accuracy. Unlike training hyperparameters, which have predictable effects, model parameters such as $\alpha$, $\beta$, and $\gamma$ directly influence model performance.
Experiments are conducted on the NTU-RGB+D XSub dataset. As shown in Table 9, different configurations of the hyperparameters are tested. The best performance, achieving 78.5% accuracy, results from a balanced configuration ($\alpha_{b}=0.25,\alpha_{j}=0.25,\alpha_{bm}=0.25,\alpha_{jm}=0.25$, $\beta_{1}=\beta_{2}=0.25$ and $\gamma_{1}=\gamma_{2}=\gamma_{3}=1.0$).
The results demonstrate that MK-SGN’s accuracy is sensitive to these model hyperparameters. Variations from the optimal configuration lead to a decrease in accuracy, highlighting the importance of tuning these parameters for optimal performance.
4.3 Visualization Analysis
Spiking Self-Attention Visualization
To elucidate how MK-SGN makes decisions and to quantify the contribution of spiking attention across depth and time, we visualize the spike self-attention maps produced by the SSA module, in Fig.5. Concretely, we render the aggregated $k\!\odot\!v$ maps (summed over heads) at each layer and along the temporal axis; higher grayscale values indicate stronger spike-modulated attention. The horizontal axis indexes the NTU 25 joints, and the vertical axis indexes time steps $T$.
(a) Spiking attention maps for Drop.
(b) Spiking attention maps for Put on Glasses.
(c) Spiking attention maps for Hopping.
Figure 5: Spiking attention maps for different actions. The x-axis represents NTU 25 joints, and the y-axis corresponds to time steps $T$. The grayscale intensity indicates the activation level of the attention mechanism at each joint over time.
Case A: Drop. In Fig. 5(a), early layers (L1–L2) exhibit diffuse activations. From L3, attention concentrates on the upper-limb chain (shoulder–elbow–wrist–hand, joints 5–12), which directly governs the dropping motion. Moderate responses on the spine chain (joints 1–3, 21) indicate posture stabilization, while the lower-limb chains (joints 13–20) remain weak. In L6, activity condenses into sparse temporal bands aligned with release and follow-through, showing SSA’s ability to emphasize task-critical skeletal substructures.
Case B: Put on Glasses. In Fig. 5(b), initial activations are diffuse. From L3, attention intensifies on the hand–arm chains (shoulder–elbow–wrist–hand, joints 5–8, 9–12) and their connection to the head chain (neck–head, joints 3–4, 21). This highlights the coordinated interaction of arms and head required to lift and place glasses. In later layers, spikes are temporally localized around approach and placement, emphasizing fine-grained motor control in upper-limb–head coordination.
Case C: Hopping. As shown in Fig. 5(c), early layers (L1–L2) present distributed activations. From L3, focus shifts toward the lower-limb chains (hip–knee–ankle–foot, joints 13–20), which dominate propulsion and landing. In deeper layers, the spine chain (joints 1–3, 21) gains attention, reflecting its stabilizing role during mid-air balance. Final layers show temporally sharp peaks near takeoff and landing, matching the energy-intensive phases of hopping.
This analysis reveals how the attention mechanism evolves to prioritize different joints depending on the action, with the model focusing more on the relevant joints at each phase of the action.
Visualization of CAMs on Temporal and Spatial Dimensions
To demonstrate the MK-SGN model’s effectiveness in capturing key features, we apply Class Activation Mapping (CAM) to the outputs of the SGC and STC layers, as shown in Fig.6. The spatial CAM highlights activations from the SGC layer, focusing on relevant joints, while the temporal CAM reflects activations from the STC layer, emphasizing critical time steps. This visualization clearly shows how the model attends to the most important body parts and action phases for accurate classification.
Figure 6: CAM visualizations on spatial and temporal dimensions. Each subfigure corresponds to one action class, with the upper skeleton plot showing CAM intensities across spatial joints and the lower bar showing temporal contributions. Warmer colors (yellow/white) denote stronger activation responses, indicating which joints and time frames are most critical for the model’s classification.
Case A: Drop. For the Drop action, the CAM in Fig.6(a) shows that MK-SGN focuses on the upper-limb joints, particularly the shoulders, elbows, and wrists, which are crucial during the release phase. The trunk joints are moderately activated, reflecting their role in posture stabilization, while the lower-limb joints receive slight attention. In the temporal domain, peak activations occur around frames 7–8, when the object is released.
Case B: Put on Glasses. For the Put on Glasses action (Fig.6(b)), the CAM highlights the face-related joints, especially the eyes, which receive the highest activation, aligning with their significance in the action. The lower limbs and trunk show little attention, reflecting their relatively low involvement. In the temporal domain, activations peak around frames 7–8, capturing the critical moment when the glasses are placed.
Case C: Hopping. For the Hopping action (Fig.6(c)), MK-SGN concentrates attention on the lower limbs, particularly the knees and ankles, which dominate the motion. The upper body receives slight attention, consistent with the nature of the action. In the temporal domain, intermittent peaks are observed, especially near frames 9–10, reflecting the recurrent critical moments of the take-off process captured by the model.
In all three cases, MK-SGN’s CAM visualizations confirm its ability to focus on the most relevant spatial and temporal features for each action. The model selectively attends to key joints and critical time steps, proving its effectiveness in capturing the necessary cues for accurate action recognition.
t-SNE Visualization of Action Embeddings.
Figure 7: t-SNE visualization of action embeddings. The plot shows the embeddings of six action samples across four variants of MK-SGN: (a) Base-SGN, (b) MK-SGN, (c) MK-SGN (+SMF), and (d) MK-SGN (+SMF+Full KD). The clusters demonstrate the effectiveness of each model variant in capturing and distinguishing actions representation.
Fig. 7 presents the t-SNE visualization of six sample actions across four MK-SGN variants: Base-SGN, MK-SGN, MK-SGN (+SMF), and MK-SGN (+SMF+Full KD).
For Base-SGN, the embeddings show scattered distributions with slight separation between actions, indicating limited feature discrimination. MK-SGN, on the other hand, demonstrates significant improvement, with clearer clusters forming around individual actions, highlighting the effectiveness of spiking graph convolution and attention mechanisms. The addition of SMF further refines the clusters, showing improved separation between actions. Finally, the Full KD variant achieves the most compact clusters, suggesting that knowledge distillation enhances the model’s ability to learn more discriminative action representations.
This analysis underscores the incremental improvements in action representation, validating the effectiveness of each component in enhancing the model’s ability to capture and distinguish complex action features.
Table 10: Model’s architecture for NTU-RGB+D
@@ 4.3 Visualization Analysis
final 256/class number GAP-FC 256/class number 256/class number SN-GAP-FC 256/class number SN-GAP-FC
4.4 Comparison with State-of-the-Art Methods
This section comprehensively compares our proposed MK-SGN and SOTA methods for skeleton-based action recognition. The comparison is divided into two parts: (1).model architecture comparison and (2).metric comparison.
In the model architecture comparison, we focus on the configuration and layer design of MK-SGN in contrast to vanilla GCN and SNN models, using the NTU-RGB+D 60 dataset as a case study. By analyzing the architectural differences, this comparison provides a foundation for the subsequent evaluation of performance metrics, ensuring a fair and consistent basis for comparison.
In the metric comparison, we evaluate MK-SGN against SOTA models across several dimensions, including accuracy, number of operations (OPs), and theoretical energy consumption. This evaluation encompasses publicly available results and those reproduced under consistent experimental settings, providing a comprehensive perspective on MK-SGN’s ability to balance high performance with energy efficiency, particularly in energy-constrained environments.
@@ 4.4.1 Model Architecture Comparison
Table 10 summarizes the architectural differences between our proposed MK-SGN and various models, including the unified framework of Vanilla GCN [20, 21, 22, 23] and Spikformer [36], as applied to the NTU-RGB+D dataset. This comparison highlights how each model is structured to handle skeleton-based action recognition tasks.
The Vanilla GCN model is a classic ANN-based approach, utilizing GC layers to capture spatial dependencies and TC layers for temporal modeling. Its architecture typically follows a pyramid design with 10 layers of GC-TC operations, featuring progressively increasing feature dimensions and temporal downsampling through strides, culminating in a GAP-FC layer for final classification.
The Base-SGN* model serves as the baseline SNN structure, leveraging SGC and STC layers to process spike-form data. It follows a similar configuration to GCN but is adapted for spiking computation.
The Spikformer model, a spiking-based Transformer, incorporates SSA layers and Multilayer Perceptron (MLP) components for feature extraction. For this study, we modified Spikformer to align with MK-SGN in terms of layer count and hidden dimensions, ensuring a fair comparison of metrics such as energy efficiency and accuracy. This adjustment enables us to evaluate the performance of Spikformer in a setting directly comparable to MK-SGN. Similarly, other network variations, such as Spiking-driven-former, Spiking-driven-former V2, and Spiking Wavelet Transformer, used in the study are adapted to maintain consistency across evaluations.
@@ 4.4.2 Metric Comparison
Table 11: Comparative results on NTU RGB+D and NTU RGB+D 120. We evaluate our model in terms of classification accuracy (%). XSub, XView and XSet represent Cross-Subject, Cross-View and Cross-Setup splits.
@@
Param OPs Power XSub XView XSub XSet
(M) (G) (mJ) (120) (120)
@@
STAtten[9]CVPR 2025 SNN Spiking Transformer 4 3.19 1.98 2.48 72.8 79.7 60.3 61.7
MK-SGN (Ours) SNN SGN 4 2.17 0.67 0.596 78.5 85.6 67.8 69.5
@@
We compare our proposed MK-SGN with recent state-of-the-art (SOTA) ANNs and SNNs in skeleton-based action recognition, focusing on two key metrics: OPs and theoretical energy consumption. The results are summarized in Table 11 and Table 12.
@@
To ensure a fair comparison, results for existing multi-modal GCN methods use the maximum number of modalities specified in their respective papers, including joint, bone, joint motion, and bone motion. The OPs and theoretical energy consumption for these multi-modal methods are calculated by multiplying the single-stream values by the number of modalities used. Additionally, we reproduce the results of recent SOTA SNNs using only the joint modality for a direct comparison with MK-SGN. These models are adapted to align with MK-SGN in terms of layers and hidden dimensions, ensuring consistency across experimental settings.
@@
The comparison, shown in Table 11, highlights MK-SGN’s exceptional balance between accuracy and energy efficiency. MK-SGN significantly reduces theoretical energy consumption, requiring only 0.596 mJ per sample, representing a 96.28% reduction compared to ST-GCN (16.01 mJ), 99.67% reduction compared to 2S-AGCN (182.87 mJ), and 99.75% reduction compared to MS-G3D (239.51 mJ). For lightweight models like Shift-GCN (46 mJ), CTR-GCN (36.25 mJ) and BlockGCN (32.57 mJ), MK-SGN reduces energy consumption by 98.70%, 98.36% and 98.75%, respectively, demonstrating its superior efficiency across ANN-based GCN models.
Beyond energy savings, MK-SGN outperforms recent SNN models in both accuracy and efficiency. For NTU-RGB+D, MK-SGN achieves 78.5% (XSub) and 85.6% (XView), surpassing competitive SNN models such as Spikformer (73.9% XSub, 80.1% XView) and Spike-driven Transformer V2 (77.4% XSub, 83.6% XView). Notably, MK-SGN achieves this with significantly reduced energy consumption—only 0.596 mJ per sample, compared to Spikformer’s 2.17 mJ and Spike-driven Transformer V2’s 2.91 mJ. Furthermore, MK-SGN demonstrates a highly efficient design, with only 2.17M parameters, far fewer than Spike-driven Transformer V2’s 11.47M. On the NTU-RGB+D 120 dataset, MK-SGN continues to outperform SNN models, achieving 67.8% (XSub) and 69.5% (XSet), while Spikformer and Spike-driven Transformer V2 achieve lower accuracies, 61.7% (XSub) and 63.7% (XSet), and 64.3% (XSub) and 65.9% (XSet), respectively. This illustrates MK-SGN’s ability to maintain competitive performance and efficiency across different datasets.
@@
On the NW-UCLA dataset, as shown in Table 12, MK-SGN achieves 92.3% accuracy, surpassing other SOTA SNN models like Spikformer (85.4%) and Spike-driven Transformer V2 (89.4%), while maintaining significantly lower energy consumption (0.207 mJ per sample), compared to Spike-driven Transformer V2 (0.914 mJ). This confirms MK-SGN’s remarkable ability to balance accuracy and energy efficiency, especially in scenarios that demand low computational cost.
To further illustrate the synergy between accuracy and energy efficiency, we present the model comparison in Figure 8. The results, based on the NTU-RGB+D CS dataset, highlight MK-SGN’s superior performance in terms of both accuracy and power consumption.
These results validate MK-SGN as a reliable and energy-efficient solution for skeleton-based action recognition, excelling in scenarios requiring high accuracy and low computational overhead. The spiking-driven architecture of MK-SGN sets a new benchmark for balancing performance and efficiency in SNNs.
Figure 8: Comparison of accuracy and power consumption for various models on the NTU-RGB+D CS dataset. MK-SGN outperforms other models in both accuracy and energy efficiency.
4.5 Discussion
Principle Analysis: Accuracy and Efficiency Trade-off
MK-SGN achieves significant energy efficiency, but its accuracy is limited by the sparsity of SNNs. Unlike ANNs, which transmit information continuously, SNNs only transmit during spikes, leading to sparse information transfer and lower entropy.
Information entropy $H(X)$ is defined as:
$H(X)=-\sum_{i=1}^{n}p_{i}\log_{2}p_{i}$ (38)
For ANNs, continuous activations result in higher entropy, enabling richer information flow. In contrast, SNNs transmit information only when neurons spike, which limits entropy and reduces the system’s capacity to capture detailed patterns[3]. Despite its efficiency, MK-SGN’s accuracy is constrained by this sparsity. It outperforms other SNN models but still lags behind ANN-based models in high-precision tasks. The trade-off between energy efficiency and accuracy remains a challenge.
Application Scenarios
While SNNs offer exceptional energy efficiency, their accuracy is often lower than that of GCNs, especially in tasks requiring detailed pattern recognition. SNNs are particularly suited for real-time applications where energy efficiency is a priority. For example, in wearable health monitoring systems, SNNs can be used for continuous gesture and activity recognition with relatively low energy consumption, making them ideal for long-term, battery-powered use. Similarly, in mobile robotics, SNNs can process sensor data efficiently, enabling real-time action recognition and decision-making on low-power devices, such as drones or robotic assistants. These applications require real-time, low-energy operation, where the trade-off between accuracy and energy efficiency is critical.
In contrast, GCNs excel in tasks requiring high accuracy and detailed spatial and temporal relationships, such as skeleton-based action recognition in complex environments. GCNs are ideal for scenarios where precise behavior analysis is necessary, such as in surveillance systems for human activity recognition in security or healthcare settings. For example, in smart home monitoring systems, GCNs can analyze detailed human motion patterns to detect falls, abnormal behaviors, or to improve user experience through activity-based control. Similarly, in sports analytics, GCNs can be used to analyze player movements in a game, providing insights into strategies, player health, and performance. These scenarios require high accuracy to capture the fine-grained details of human motion, which GCNs are well-suited for.
Future Work
To enhance the performance of MK-SGN, future work should focus on optimizing spike coding, increasing spike rates, or exploring hybrid architectures with ANNs, like GCNs or CNNs, to enhance both accuracy and efficiency. SNNs can benefit from advanced spike encoding schemes that allow for richer information transmission while maintaining low energy consumption. Hybrid models that combine the strengths of SNNs and GCNs can offer a balance between real-time energy efficiency and high accuracy, especially in complex action recognition tasks.
Additionally, training methods tailored to SNNs should be developed to improve performance in high-accuracy tasks. Approaches such as spike-timing dependent plasticity (STDP) or supervised learning algorithms specifically designed for spiking neurons could lead to better convergence and higher recognition accuracy in the context of skeleton-based action recognition. These advancements will help address the current limitations and broaden the application of SNN-based models in real-world scenarios.
@@ 5 Conclusion
In this work, we propose the Spiking Graph Convolutional Network (MK-SGN), a novel architecture that combines the strengths of SNNs and GCNs for skeleton-based action recognition. To the best of our knowledge, this framework is the first to effectively integrate spiking neural computation with graph-based modeling, offering a new perspective on energy-efficient and high-performance recognition methods.
The MK-SGN addresses the critical challenge of high energy consumption in traditional GCN-based methods while maintaining competitive performance. Leveraging multimodal fusion and knowledge distillation, MK-SGN optimizes spiking neural computation and enhances the ability to extract and integrate meaningful features across modalities. These innovations enable MK-SGN to substantially decrease energy consumption across three benchmark datasets for skeleton action recognition, achieving an impressive reduction of over 98% compared to the most effective ANN-based methods.
Moreover, MK-SGN consistently outperforms SOTA SNN models in recognition accuracy, achieving an average improvement of 5%. This demonstrates its ability to balance energy consumption with recognition performance, making it well-suited for resource-constrained applications.However, the accuracy of MK-SGN is still limited by the sparsity inherent in SNNs. While the model achieves significant energy efficiency, its performance in high-precision tasks still lags behind traditional ANN-based methods, primarily due to the sparse information transfer characteristic of spiking networks.
Future work should focus on addressing these limitations by exploring advanced spike coding techniques to improve information transmission and enhance accuracy. Additionally, incorporating hybrid models that combine the strengths of SNNs and ANNs could help balance energy efficiency and accuracy more effectively. Optimizing spike rates and developing tailored training methods for SNNs, could also boost performance in high-accuracy scenarios.
Despite these challenges, MK-SGN offers a promising approach for energy-efficient skeleton-based action recognition, and its potential for real-time applications in resource-constrained environments positions it as a solid foundation for future exploration in spiking-based multimodal frameworks.
6 Acknowledgements
The National Natural Science Foundation of China NSFC-61976022 supports this work.